1.      The following data refers to the price of a good ‘P’ and the quantity of the good supplied, ‘S’.

p 2 7 5 1 4 8 2 8

___/_____________________________________

s / 15 41 32 9 24 43 17 40

a.      Estimate the linear regression line

b.     Estimate the standard errors of

c.      Test the hypothesis that price influences supply



1
Expert's answer
2021-07-08T15:20:34-0400

The linear regression line is "S=\alpha+\beta P+E"

E is random error

and "E(s)=\alpha+\beta P"

"S_{SS}=\sum S^2=\displaystyle\sum_{i=1}^8(S_i-S)^2=1205"

"S_{PP}=\sum P^2= \displaystyle\sum_{i=1}^8(P_i-P)^2=55.9"

"S_{SP}=\sum(SP)=\displaystyle\sum_{i=1}^8(S_i-S)(P_i-P)=22.4"

"\alpha=S-P\beta" and "\beta=\frac{(\sum SP)}{(\sum S^2)^{1/2}\times (\sum P^2)^{1/2}}" "=\frac{(225.4)}{(1205)^{1/2}\times (55.9)^{1/2}}=0.8685"

from the table "S=\sum S_i/n=225/8=28.125"

"P=\sum P_i/n=37/8=4.625"

"\alpha=28.125-4.625\times0.8685=24.1082"

a) the estimated regression line is, "S=24.1082+0.8685P_i"


b) the standard error (SE) of "\alpha" and "\beta" are

"SE(\alpha)=\sigma \sqrt{1/n+P^2/S_{pp}}"

and "SE(\alpha)=\sigma/\sqrt{S_{pp}}"

"\sigma^2=1/(n-2)SSE=1/(n-2)[S_{ss}-\beta^2S_{pp}]=1/(8-2)[1205-0.8685^2\times 55.9]"

"=1/6\times1162.8351=193.8058"

"\sigma=\sqrt{193.8058}=13.9214"

"SE(\alpha)=\sigma \sqrt{1/n+P^2/S_{pp}}=13.9214/\sqrt{55.9}=13.9214/7.4766=1.86199"


c) testing for hypothesis 

"H_0:\beta=0 \space versus \space H_1:\beta \not= 0"

at "\alpha=0.05"

"t=(\beta-0)/(SE(\beta))"

"t=0.8685\/1.86199=0.4664"

"t_{critical}=2.4469"

"|t|\lt2.4469"

we fail to reject "H_0" , that means at "\alpha=0.05, P\space doesn't\space affect \space S"